# Sourced Carrollian Fluids Dual to Black Hole Horizons

### v2

@@ V.2 Numerical Results
We solve the horizon equations recast in the spin-weighted formalism with a fourth-order Runge–Kutta integrator in Python. [^9] We make extensive use of the Python libraries NumPy [78], Matplotlib [79], SymPy [80], and SciPy [81]. We obtain the time evolution of the spin-weighted scalars $\{\vartheta,H,C,K\}$, from which we recover the respective geometric quantities on the horizon, $\{\Theta,\mathcal{H}_{A},q_{AB},\kappa\}$.
The system of equations at hand consists of four coupled ordinary differential equations. Due to the teleological nature of the event horizon, the system is set up as a Schwarzschild black hole at a large null time $v$, and the boundary condition that the perturbations diminish at large $v$ is imposed. Then, the equations are solved from the future to the past of the evolution. (See [65] for more details of “backward-in-time” evolution.) Therefore, the boundary conditions imposed at the distant future of the horizon are that the null expansion and the Hájíček field vanish at all perturbative orders. At the same time, the 2-metric becomes the metric on the unit sphere, and the surface gravity assumes its value for a static Schwarzschild black hole.
@@ V.2.1 Quadratic quasinormal modes
Quadratic quasinormal modes have been of interest for the study of nonlinear effects in gravity [70, 71, 73, 72]. Here, we focus on the generation of quadratic modes from the linear modes of source fields on the black hole event horizon.
QNM frequencies of the scalar field source modes combine quadratically through the horizon equations. This follows from the fact that the stress-energy tensor is quadratic in the scalar field, so are the corresponding source terms, $S_{i}\propto(\partial_{v}\Phi\partial_{i}\Phi)$. Therefore, we expect these quadratic QNM frequencies to appear in the evolution of the geometric quantities. We extract the frequencies from our result using the QNM rational filter [82, 83, 84]. These results are summarized in Fig. 3.
@@
To further confirm that these modes are indeed quadratic QNMs, we check the amplitudes as we change the amplitude of the scalar field profile. For example, we observe that the amplitude of $\vartheta_{40}$ scales quadratically with the amplitudes of $\varphi_{20}$, further supporting that these modes are generated as the quadratic combinations of linear modes.
@@ V.2.2 Horizon area
On the other hand, the energy flux due to the scalar field passing through the horizon is given by [85]
@@
Since the perturbation parameter $\epsilon$ is small, the change in the black hole parameters through the entire time evolution can be assumed quasistatic. In this regime, one can relate these two expressions via the first law of black hole mechanics [^10] See [86] for a review on black hole thermodynamics.
@@ V.2.3 Horizon angular momentum
The flux of angular momentum through the event horizon is calculated as follows [85]:
@@
Assuming that the deviation from axisymmetry is small through the perturbative evolution, the axial Killing vector in the perturbed spacetime is expressed by $\psi^{a}\equiv\partial_{\phi}$. Note that these expressions hold within the assumption that the perturbed black hole deviates slightly from the Kerr solution as discussed in [85]. In the perturbative expansion, the above expression becomes
@@
On the other hand, the DNS equation can be used to relate the Hájíček field to the angular momentum on the event horizon through Damour’s original prescription [3]
@@
Contrarily, the angular momentum computed from the Hájíček field has a completely different behavior. The solid orange curve rises exponentially before $v=0$. This is due to the secular terms in Eq.s 75 & 76, which permeate to the angular momentum computed from Eq. 93. Therefore, at some $v$ around $v=0$, the perturbative approximation fails to capture the physical behavior.
@@ V.2.5 Vorticity of $\mathcal{H}_{A}$
Finally, we compute the vorticity of the Hájíček field defined by [55]
@@ VI Conclusions
We plan to extend our results to tackle the Carrollian fluid dual to a black hole horizon in a full nonlinear treatment. This could include a dynamical approach to hydrodynamical horizons [57], which is briefly discussed at the end of Appendix B.2. Another natural extension of this work is to build on the discussions in [55] to study the nonlinear-to-linear transition and explore enstrophy definitions within Carrollian fluids and the counterpart in gravity, if such a concept exists for such fluids.
A formalism similar to this work can be used to investigate the Carrollian fluid dual to any other null surface at null infinity [87] or null surfaces in the bulk not corresponding to an event horizon. Notice that, in such a case, the surface will always expand (if outside the horizon) or always contract (if inside of it), and thereby the associated Carrollian energy will not equilibrate. Finally, on the black hole perturbation front, the horizon equations we studied in this work can yield insights about the dynamics of the horizon, and its relation to quasinormal mode behavior and the non-linear/linear transition (See, for example, [67] ).
@@ Acknowledgements.
We would like to thank Luca Ciambelli, Puttarak Jai-akson, Jaime Redondo-Yuste, and Sizheng Ma for useful discussions. LL acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery Grant. LL also thanks financial support via the Carlo Fidani Rainer Weiss Chair at Perimeter Institute and CIFAR. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ Appendix A Derivatives & connections
The prescription [57] we followed to construct a null hypersurface involved projecting quantities in the bulk onto different hypersurfaces, as discussed in the main text. Here, we summarize how the covariant derivatives and curvature connections are induced in such hypersurfaces.
The rigged derivative $D_{i}$ is defined [56] as the projection of the covariant derivative $\nabla_{a}$ on $\mathscr{M}$ onto $\mathscr{H}$. For a tensor with mixed indices on $\mathscr{H}$, the rigged derivative acts as the following
@@
where it is useful to note $q_{A}^{\hskip 4.0pta}=\left(e_{A}\right)^{a}$. This derivative satisfies $\hat{\nabla}_{C}\,q_{AB}=0$ and preserves Carrollian diffeomorphism; therefore, it is often referred to as the Levi-Civita–Carroll covariant derivative [32]. Related to this covariant derivative, a Levi-Civita–Carroll connection is introduced. The Levi-Civita–Carroll covariant derivative acts on a vector as
@@
where $\hat{\partial}_{A}=\partial_{A}+\frac{b_{A}}{\alpha}\partial_{t}$, and $\hat{\gamma}^{B}_{AC}$ is the Levi-Civita–Carroll connection [32].
The commutation of the Levi-Civita–Carroll covariant derivative gives rise to the Riemann–Carroll tensor, which is covariant under Carrollian diffeomorphisms [32]:
@@
The Ricci–Carroll tensor can be expressed in terms of the geometric quantities defined on the horizon as follows [57]:
@@
Note that, within the gauge expressed in the main text, yielding the metric in Eq. (II.2), the following choices are made: $\alpha=1$ and $b_{A}=0$. Therefore, the structures introduced in this appendix simplify accordingly. Specifically, the Riemann–Carroll and Ricci–Carroll tensors agree with the induced Riemann and Ricci tensors on the hypersurfaces. See the Appendix sections of [32] and [57] for more details on derivatives and connections summarized here.
@@ Appendix B Gauge choices on the stretched Carrollian formalism
Starting from the general metric on the bulk spacetime given in section 3 of [57], we imposed some conditions as discussed in the main text. Here, we summarize other possible choices and their implications for the horizon and Carrollian dynamics.
@@ B.2 Choices on $b_{A}$ and $V_{A}$
To construct the metric (Eq. (II.2)), we used $b_{A}=0$ and $V_{A}\neq 0$. Instead, if we had imposed $V_{A}=0$ and $b_{A}\neq 0$, then we would have a non-vanishing Carrollian acceleration $\varphi_{A}$, vorticity $\omega_{AB}$ on the horizon alongside a nonzero Hájíček field. The Carrollian acceleration explicitly appears in the horizon equations under this coordinate choice. Moreover, the covariant derivative in the expressions has to be changed with the Carroll covariant derivative that preserves Carrollian diffeomorphisms [32] as summarized in Appendix A. The associated partial derivative acts on a scalar function $f\left(x^{i}\right)$ on $\mathscr{H}$ as
@@
If instead, we follow the conventions of Randers–Papapetrou gauge in [21, 32, 56] to have $b_{A}{\left(x^{i}\right)}\neq 0$ and $V_{A}=0$, while still keeping $\alpha=1$, we make the following intricate observation regarding the dynamics of the set of projected Einstein field equation on the horizon. We find that both $\varphi_{A}$ and $\mathcal{H}_{A}$ depend on $b_{A}$ on the horizon as
@@
Hence, the extra term of $\varphi_{A}$ appearing in the DNS (Eq. (28)) is not an independent function; rather, it satisfies $\varphi_{A}=-2\mathcal{H}_{A}$. This, however, changes the character of the time evolution of the Hájíček field. The exponential behavior of the field converges to 0 as the null time evolves to infinity. This behavior is the opposite of that of the null expansion, which diverges in the infinite future as the Raychaudhuri equation (Eq.  (27)) dictates. This well-known “problem” in the horizon evolution is explained by its teleological nature [4].
@@
The teleological nature of the event horizon is also the reason why we implement a backward-in-time evolution to solve the horizon equations [65].
@@
The general case as studied in [57] sets both $b_{A}\neq 0$ and $V_{A}\neq 0$. Then the Hájíček field near the horizon behaves as
@@
This, however, introduces $b_{A}$ as an extra degree of freedom in the horizon equations. Therefore, more constraints need to be satisfied to solve the system of equations self-consistently. A set of such constraints called hydrodynamical horizon is studied in [57]. We intend to approach the resulting system with the numerical methods developed in this paper in future work.
@@ Appendix C Horizon equations in the perturbative scheme
Here, we explicitly express the horizon equations expanded in the perturbative treatment up to $\mathcal{O}\left(\epsilon^{2}\right)$.
The equations in order $\mathcal{O}\left(1\right)$, that is, the order $\epsilon^{0}$, are identically satisfied by the background Schwarzchild metric.
@@
    $\partial_{v}\mathcal{H}_{A}^{(1)}-{\nabla}_{A}\kappa^{(1)}+\Omega^{\text{(Sph)}{BC}}{\nabla}_{B}\sigma_{AC}^{(1)}=0$ (115)
    $\kappa^{(1)}+M{\nabla}_{A}\mathcal{H}^{A(1)}-\frac{M}{2}\hat{R}^{(1)}=0$ (116)
    $2\partial_{v}{\bar{\sigma}}_{AB}^{(1)}+\frac{1}{2M}\bar{{\sigma}}_{AB}^{(1)}+\frac{1}{M}\sigma_{AB}^{(1)}+2{\nabla}_{\langle A}\mathcal{H}_{B\rangle}^{(1)}=0$ (117)
@@
    $\partial_{v}\Theta^{(2)}-\frac{1}{4M}\Theta^{(2)}+\sigma^{AB(1)}\sigma_{AB}^{(1)}=-\left(\partial_{v}\Phi^{(1)}\right)^{2}$ (118)
    $\partial_{v}\mathcal{H}_{A}^{(2)}+{\nabla}_{B}\left(\Omega^{\text{(Sph)}{BC}}\sigma_{AC}^{(2)}\right)-{\nabla}_{A}\kappa^{(2)}-\frac{1}{2}{\nabla}_{A}\Theta^{(2)}=\partial_{v}\Phi^{(1)}\partial_{A}\Phi^{(1)}$ (119)
    $\partial_{v}{\bar{\Theta}}^{(2)}+\frac{1}{4M}{\bar{\Theta}}^{(2)}+\frac{1}{M}\left(\kappa^{(2)}+\Theta^{(2)}\right)+{\nabla}_{A}\mathcal{H}^{A(2)}+\mathcal{H}^{(1)}_{A}\mathcal{H}^{A(1)}-\frac{1}{2}\hat{R}^{(2)}=-\Omega^{\text{(Sph)}{AB}}\partial_{A}\Phi^{(1)}\partial_{B}\Phi^{(1)}$ (120)
    $2\partial_{v}{\bar{\sigma}}_{AB}^{(2)}-2\Omega^{\text{(Sph)}}_{AB}\left(\bar{\sigma}_{CD}^{(1)}\sigma^{CD(1)}\right)+\frac{1}{2M}\bar{{\sigma}}_{AB}^{(2)}+2\kappa^{(1)}\bar{{\sigma}}_{AB}^{(1)}+\frac{1}{M}\sigma_{AB}^{(2)}+2{\nabla}_{\langle A}\mathcal{H}_{B\rangle}^{(2)}+2\mathcal{H}^{(1)}_{\langle A}\mathcal{H}^{(1)}_{B\rangle}$
    $\hskip 199.16928pt=-{\partial}_{A}\Phi^{(1)}{\partial}_{B}\Phi^{(1)}+\frac{1}{2}\Omega^{\text{(Sph)}}_{AB}\left(\Omega^{\text{(Sph)}CD}{\partial}_{C}\Phi^{(1)}{\partial}_{D}\Phi^{(1)}\right)$ (121)

### v3

@@ V.2 Numerical Results
We solve the horizon equations recast in the spin-weighted formalism with a fourth-order Runge–Kutta integrator in Python. [^12] We make extensive use of the Python libraries NumPy [harris2020array], Matplotlib [Hunter:2007], SymPy [10.7717/peerj-cs.103], and SciPy [2020SciPy-NMeth]. We obtain the time evolution of the spin-weighted scalars $\{\vartheta,H,C,K\}$, from which we recover the respective geometric quantities on the horizon, $\{\Theta,\mathcal{H}_{A},q_{AB},\kappa\}$.
The system of equations at hand consists of four coupled ordinary differential equations. Due to the teleological nature of the event horizon, the system is set up as a Schwarzschild black hole at a large null time $v$, and the boundary condition that the perturbations diminish at large $v$ is imposed. Then, the equations are solved from the future to the past of the evolution. (See [G_mez_2001] for more details of “backward-in-time” evolution.) Therefore, the boundary conditions imposed at the distant future of the horizon are that the null expansion and the Hájíček field vanish at all perturbative orders. At the same time, the 2-metric becomes the metric on the unit sphere, and the surface gravity assumes its value for a static Schwarzschild black hole.
@@ V.2.1 Quadratic quasinormal modes
Quadratic quasinormal modes have been of interest for the study of nonlinear effects in gravity [Ma_2024, Bucciotti_2024, Khera_2025, Khera_2023]. Here, we focus on the generation of quadratic modes from the linear modes of source fields on the black hole event horizon.
QNM frequencies of the scalar field source modes combine quadratically through the horizon equations. This follows from the fact that the stress-energy tensor is quadratic in the scalar field, so are the corresponding source terms, $S_{i}\propto(\partial_{v}\Phi\partial_{i}\Phi)$. Therefore, we expect these quadratic QNM frequencies to appear in the evolution of the geometric quantities. We extract the frequencies from our result using the QNM rational filter [Ma_2022, Ma_2023, Ma_2023_2]. These results are summarized in Fig. 3.
@@
To further confirm that these modes are indeed quadratic QNMs, we check the amplitudes as we change the amplitude of the scalar field profile. For example, we observe that the amplitude of $\vartheta_{40}$ scales quadratically with the amplitudes of $\varphi_{20}$, further supporting that these modes are generated as quadratic combinations of linear modes.
@@ V.2.2 Horizon area
On the other hand, the energy flux due to the scalar field passing through the horizon is given by [Poisson_2004]
@@
Since the perturbation parameter $\epsilon$ is small, the change in the black hole parameters through the entire time evolution can be assumed quasistatic. In this regime, one can relate these two expressions via the first law of black hole mechanics [^13] See [Wald_2001] for a review on black hole thermodynamics.
@@ V.2.3 Horizon angular momentum
The flux of angular momentum through the event horizon is calculated as follows [Poisson_2004]:
@@
Assuming that the deviation from axisymmetry is small through the perturbative evolution, the axial Killing vector in the perturbed spacetime is expressed by $\psi^{a}\equiv\partial_{\phi}$. Note that these expressions hold within the assumption that the perturbed black hole deviates slightly from the Kerr solution as discussed in [Poisson_2004]. In the perturbative expansion, the above expression becomes
@@
On the other hand, the DNS equation can be used to relate the Hájíček field to the angular momentum on the event horizon through Damour’s original prescription [Damour1982]
@@
Contrarily, the angular momentum computed from the Hájíček field has a completely different behavior. The solid orange curve rises exponentially before $v=0$. This is due to the secular terms in Eq.s 70 & 71, which permeate to the angular momentum computed from Eq. 88. Therefore, at some $v$ around $v=0$, the perturbative approximation fails to capture the physical behavior.
@@ V.2.5 Vorticity of $\mathcal{H}_{A}$
Finally, we compute the vorticity of the Hájíček field defined by [Redondo_Yuste_2023]
@@ VI Conclusions
The full equations describing horizon dynamics are nonlinear. Therefore, one expects that the Carrollian fluid/horizon geometry duality should extend to the general case beyond a perturbative treatment. Nevertheless, we observe the hints of nonlinear effects, such as the generation of quadratic QNMs on the horizon, which can provide us with intuition for the nonlinear effects within fluid/gravity correspondence. Resolving the implications of this duality in a full dynamical setting would require a more sophisticated numerical approach and is left for future work. This could include a dynamical approach to hydrodynamical horizons [freidel2024geometrycarrollianstretchedhorizons], which is briefly discussed at the end of Appendix B.2. Another natural extension of this work is to build on the discussions in [Redondo_Yuste_2023] to study the nonlinear-to-linear transition and explore enstrophy definitions within Carrollian fluids and the counterpart in gravity, if such a concept exists for such fluids.
A formalism similar to this work can be used to investigate the Carrollian fluid dual to any other null surface at null infinity [Ciambelli:2025mex] or null surfaces in the bulk not corresponding to an event horizon. Notice that, in such a case, the surface will always expand (if outside the horizon) or always contract (if inside of it), and thereby the associated Carrollian energy will not equilibrate. Finally, on the black hole perturbation front, the horizon equations we studied in this work can yield insights about the dynamics of the horizon and its relation to quasinormal mode behavior and the nonlinear-to-linear transition (see, for example, [May_2024]).
@@ Acknowledgements.
We thank Luca Ciambelli, Puttarak Jai-akson, Jaime Redondo-Yuste, and Sizheng Ma for useful discussions. LL acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery Grant. LL also thanks financial support via the Carlo Fidani Rainer Weiss Chair at Perimeter Institute and CIFAR. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
Data availability
Our code is publicly available [CodeOnGit].
Appendices
Some details of the computations and the arguments of the main text are included in the appendices. In Appendix A, we summarize the derivatives and connections relevant to the stretched Carrollian structure. In Appendix B, we expand on the gauge choices implemented in Section II. We leave some expressions in the derivations of perturbative equations to Appendix C, while expounding the numerical implementation in Appendix D. Finally, in Appendix E, we explicitly treat two well-known spacetimes in the structure introduced in Section II as concrete examples.
@@ Appendix A Derivatives & connections
The prescription [freidel2024geometrycarrollianstretchedhorizons] we followed to construct a null hypersurface involved projecting quantities in the bulk onto different hypersurfaces, as discussed in the main text. Here, we summarize how the covariant derivatives and curvature connections are induced in such hypersurfaces.
The rigged derivative $D_{i}$ is defined [freidel2022carrollian] as the projection of the covariant derivative $\nabla_{a}$ on $\mathscr{M}$ onto $\mathscr{H}$. For a tensor with mixed indices on $\mathscr{H}$, the rigged derivative acts as the following
@@
where it is useful to note $q_{A}^{\hskip 4.0pta}=\left(e_{A}\right)^{a}$. This derivative satisfies $\hat{\nabla}_{C}\,q_{AB}=0$ and preserves Carrollian diffeomorphism; therefore, it is often referred to as the Levi-Civita–Carroll covariant derivative [Ciambelli_2018]. Related to this covariant derivative, a Levi-Civita–Carroll connection is introduced. The Levi-Civita–Carroll covariant derivative acts on a vector as
@@
where $\hat{\partial}_{A}=\partial_{A}+\frac{b_{A}}{\alpha}\partial_{t}$, and $\hat{\gamma}^{B}_{AC}$ is the Levi-Civita–Carroll connection [Ciambelli_2018].
The commutation of the Levi-Civita–Carroll covariant derivative gives rise to the Riemann–Carroll tensor, which is covariant under Carrollian diffeomorphisms [Ciambelli_2018]:
@@
The Ricci–Carroll tensor can be expressed in terms of the geometric quantities defined on the horizon as follows [freidel2024geometrycarrollianstretchedhorizons]:
@@
Note that, within the gauge expressed in the main text, resulting in the metric in Eq. (II.2), the following choices are made: $\alpha=1$ and $b_{A}=0$. Therefore, the structures introduced in this appendix simplify accordingly. Specifically, the Riemann–Carroll and Ricci–Carroll tensors agree with the induced Riemann and Ricci tensors on the hypersurfaces. See the Appendix sections of [Ciambelli_2018] and [freidel2024geometrycarrollianstretchedhorizons] for more details on derivatives and connections summarized here.
@@ Appendix B Gauge choices on the stretched Carrollian formalism
Starting from the general metric on the bulk spacetime given in section 3 of [freidel2024geometrycarrollianstretchedhorizons], we imposed some conditions as discussed in the main text. Here, we summarize other possible choices and their implications for the horizon and Carrollian dynamics.
@@ B.2 Choices on $b_{A}$ and $V_{A}$
To construct the metric (Eq. (II.2)), we used $b_{A}=0$ and $V_{A}\neq 0$. Instead, if we had imposed $V_{A}=0$ and $b_{A}\neq 0$, then we would have a non-vanishing Carrollian acceleration $\varphi_{A}$, vorticity $\omega_{AB}$ on the horizon alongside a nonzero Hájíček field. The Carrollian acceleration explicitly appears in the horizon equations under this coordinate choice. Moreover, the covariant derivative in the expressions has to be changed with the Carroll covariant derivative that preserves Carrollian diffeomorphisms [Ciambelli_2018] as summarized in Appendix A. The associated partial derivative acts on a scalar function $f\left(x^{i}\right)$ on $\mathscr{H}$ as
@@
If instead, we follow the conventions of Randers–Papapetrou gauge in [Donnay_2019, Ciambelli_2018, freidel2022carrollian] to have $b_{A}{\left(x^{i}\right)}\neq 0$ and $V_{A}=0$, while still keeping $\alpha=1$, we make the following intricate observation regarding the dynamics of the set of projected Einstein field equation on the horizon. We find that both $\varphi_{A}$ and $\mathcal{H}_{A}$ depend on $b_{A}$ on the horizon as
@@
Hence, the extra term of $\varphi_{A}$ appearing in the DNS (Eq. (30)) is not an independent function; rather, it satisfies $\varphi_{A}=-2\mathcal{H}_{A}$. This, however, changes the character of the time evolution of the Hájíček field. The exponential behavior of the field converges to 0 as the null time evolves to infinity. This behavior is the opposite of that of the null expansion, which diverges in the infinite future as the Raychaudhuri equation (Eq.  (29)) dictates. This well-known “problem” in the horizon evolution is explained by its teleological nature [TPM1986].
@@
The teleological nature of the event horizon is also the reason why we implement a backward-in-time evolution to solve the horizon equations [G_mez_2001].
@@
The general case as studied in [freidel2024geometrycarrollianstretchedhorizons] sets both $b_{A}\neq 0$ and $V_{A}\neq 0$. Then the Hájíček field near the horizon behaves as
@@
This, however, introduces $b_{A}$ as an extra degree of freedom in the horizon equations. Therefore, more constraints need to be satisfied to solve the system of equations self-consistently. A set of such constraints called hydrodynamical horizon is studied in [freidel2024geometrycarrollianstretchedhorizons]. We intend to approach the resulting system with the numerical methods developed in this paper in future work.
@@ Appendix C Horizon equations in the perturbative scheme
Here, we explicitly express the horizon equations expanded in the perturbative treatment up to $\mathcal{O}\left(\epsilon^{2}\right)$. The equations in order $\mathcal{O}\left(1\right)$, that is, the order $\epsilon^{0}$, are identically satisfied by the background Schwarzchild metric.
@@
    $\partial_{v}\Omega^{(1)}_{AB}$ $=2M\partial_{v}\lambda_{AB}^{(1)}+\frac{1}{2}\lambda_{AB}^{(1)}-\frac{1}{2M}\Omega_{AB}^{(1)}-2M{\nabla}_{\langle A}\mathcal{H}_{B\rangle}^{(1)}$ (110)
    $\kappa^{(1)}$ $=-M{\nabla}_{A}\mathcal{H}^{A(1)}+\frac{M}{2}\hat{R}^{(1)}$ (112)